OpenStudy (anonymous):

Can some one help me with this problem please...I am stressing and about in tears... 13- (2c+2)=2(c+2)+3c... Please show me the steps to how you came up the answer, I need to understand..please

OpenStudy (anonymous):

13-(2c+2)=2(c+2)+3c 13-2c-2=2c+4+3c 13-2-4=2c+2c+3c 7=7c 1=c

OpenStudy (anonymous):

thanks